Your math tests scores are 68 78 90 and 91. What is the lowest grade you can get on the next test to still have an average at least 85
Answer by tinbar(133)

(68+78+90+91+lowestgrade)/5=85
now we jus need to solve backwards for the lowestgrade
(68+78+90+91+lowestgrade)=85*5
327+lowestgrade = 425
lowestgrade = 425-327
lowestgrade = 98
therefore u need AT LEAST 98